OpenStudy (anonymous):

Find the area of a polygon with vertices at (1, 3), (7, 3), (7, 7), and (4, 7).

OpenStudy (rajee_sam):

 The given polygon is a Trapezoid. If you plot the points you see (1,3) and (7,3) form a straight line parallel to x-axis and similarly (4,7) and (7,7) form a line parallel to x -axis. So in turn both these lines are parallel to each other. Now use the formula for finding the areal of a trapezoid. \[Area = \frac{ 1 }{2 } \times h \times (b _{1}+ b _{2})\] where 'h' is the vertical height and b1 and b2 are the parallel sides. Plug your values in the formula and find the area.